Find the slope from the standard form of the equation.
3x - y = 7

Answer by Alan3354(69443)

Same problem, different numbers.
=================
Find the slope of the line.
3x - 2y = 10
-----------
Change to slope-intercept form, y = mx + b
m is the slope.
----
That means solve for y:
3x - 2y = 10
2y = 3x - 10
y = (3/2)x - 5
Slope m = 3/2